# Centered Moments of Weighted One-Level Densities of $GL(2)$ $L$-Functions

### v3

@@ 5. Combinatorial Analysis
Let $\chi$ be a nontrivial character of modulus $D\geq 1$. Let $n$ be a positive integer and $m$ be a nonnegative integer such that $m\leq n$ and $m\equiv n\pmod{2}$. Then
$\sum_{\begin{subarray}{c}q\text{ prime};\\ q\nmid N\end{subarray}}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n}\frac{\chi(q)^{m}\log^{n}q}{q^{(m+n)/2}\log^{n}Q_{k,N}}\ \sigma_{1}\left(\left(r,q^{m}\right)\right)$
$=$ $\ \begin{cases}\frac{\sigma_{\phi}^{2}}{4}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)&(m,n)=(0,2),\\ O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)&(m,n)=(1,1),\\ O\left(\frac{1}{\log^{2}Q_{k,N}}\right)&\text{otherwise}.\end{cases}$ (5.9)
@@
First, we consider $(m,n)=(0,2)$. In this case, we apply Abel’s summation formula in tandem with the prime number theorem, recalling that $\phi$ is even and $\widehat{\phi}$ has compact support. Therefore, the sum under consideration equals
$\sum_{\begin{subarray}{c}q\text{ prime};\\ q\nmid N\end{subarray}}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{2}\frac{\log^{2}q}{q\log^{2}Q_{k,N}}\ =\ \frac{\sigma_{\phi}^{2}}{4}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right).$ (5.10)
@@
Second, we consider $(m,n)=(1,1)$. We follow the same method from the previous case and appeal to Dirichlet’s theorem on primes in arithmetic progressions (the value of $\chi$ is $\pm 1$ on exactly half the primes, upto lower order terms). This gives
$\sum_{\begin{subarray}{c}q\text{ prime};\\ q\nmid N\end{subarray}}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)\frac{\chi(q)\log q}{q\log Q_{k,N}}\ =\ \sum_{q\text{ prime}}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)\frac{\chi(q)\log q}{q\log Q_{k,N}}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)$
$=$ $\ \sum_{q:\chi(p)=1}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)\frac{\log q}{q\log Q_{k,N}}-\sum_{q:\chi(p)=-1}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)\frac{\log q}{q\log Q_{k,N}}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)$
$=$ $\ O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right);$ (5.11)
this equation is also given in [KR19, Page 12].
Finally, we consider $(m,n)\neq(0,2),(1,1)$. Since $m$ and $n$ are integers with the same parity, we know that $m+n\geq 4$ in this case, and the sum on the right-hand side of (5) converges:
$\sum_{\begin{subarray}{c}q\text{ prime};\\ q\nmid N\end{subarray}}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n}\frac{\chi(q)^{m}\log^{n}q}{q^{(m+n)/2}\log^{n}Q_{k,N}}\ \sigma_{1}\left(\gcd\left(r,q^{m}\right)\right)\ \ll_{r}$ $\ \frac{1}{\log^{n}Q_{k,N}}\sum_{q\text{ prime}}\frac{\log^{n}q}{q^{(m+n)/2}}$
$=$ $\ O\left(\frac{1}{\log^{2}Q_{k,N}}\right).$ (5.12)
This completes the analysis for $\chi$ nontrivial. ∎
@@
Let $\chi_{0}$ be the trivial character of modulus $D\geq 1$. Let $n$ be a positive integer and $m$ be a nonnegative integer such that $m\leq n$ and $m\equiv n\pmod{2}$. Then
$\sum_{\begin{subarray}{c}q\text{ prime};\\ q\nmid N\end{subarray}}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n}\frac{\chi_{0}(q)^{m}\log^{n}q}{q^{(m+n)/2}\log^{n}Q_{k,N}}\ \sigma_{1}\left(\gcd\left(r,q^{m}\right)\right)$
$=$ $\ \begin{cases}\frac{\sigma_{\phi}^{2}}{4}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)&(m,n)=(0,2),\\ \frac{\phi(0)}{2}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)&(m,n)=(1,1),\\ O\left(\frac{1}{\log^{2}Q_{k,N}}\right)&\text{otherwise}.\end{cases}$ (5.13)
@@
The analysis for $(m,n)\neq(1,1)$ is identical to Lemma 5.2, so we need only tend to $(m,n)=(1,1)$. We compute the corresponding sum using the methods from the previous lemma:
$\sum_{\begin{subarray}{c}q\text{ prime};\\ q\nmid N\end{subarray}}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)\frac{\chi_{0}(q)\log q}{q\log Q_{k,N}}\ =$ $\ \sum_{q\text{ prime}}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)\frac{\log q}{q\log Q_{k,N}}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)$
$=$ $\ \frac{\phi(0)}{2}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right).$ (5.14)
This equation is also given in [KR19, Page 11]. ∎
Considering these lemmas, we calculate the $n$th centered moment separately for non-trivial and trivial $\chi$ in Subsections 5.1 and 5.2, respectively. For clarity of the exposition, we let
$T(q,n_{j},m_{j},\chi)\ :=\ \widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n_{j}}\frac{\chi(q)^{m_{j}}\log^{n}q}{q^{(n+m)/2}\log^{n}Q_{k,N}}\ \sigma_{1}\left(\gcd\left(r,q^{m}\right)\right)$ (5.15)
going forward.
@@ 5.1. Non-Trivial Twisting Character
In view of Lemma 5.2, we split our consideration across two cases: $(m_{j},n_{j})\neq(0,2)$ for some $j$, and $(m_{j},n_{j})=(0,2)$ for all $j$.
In the former case, consider the $j$ for which $(m_{j},n_{j})\neq(0,2)$. For this $j$, Lemma 5.2 yields
$\sum_{q\nmid N}T(q,n_{j},m_{j},\chi)\ =\ O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right).$ (5.16)
For all the other $j$, Lemma 5.2 gives
$\sum_{q\nmid N}T(q,n_{j},m_{j},\chi)\ =\ O(1).$ (5.17)
Multiplying the order of magnitude over all $j$, we find that the main term does not contribute in this case, asymptotically speaking:
$\sum_{\begin{subarray}{c}1\leq\ell\leq t;\\ n_{1}+\dots+n_{\ell}=t;\\ m_{j}\equiv n_{j}(2)\end{subarray}}\frac{t!}{\ell!n_{1}!\cdots n_{\ell}!}\left(\prod_{j=1}^{\ell}c_{n_{j},m_{j}}\right)\prod_{j=1}^{\ell}\left(\sum_{q\nmid N}T(q,n_{j},m_{j},\chi)\right)=O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right);$ (5.18)
we absorb the combinatorial sum into the constant as it is bounded by a polynomial in $t\leq n$.
The latter case, $(m_{j},n_{j})=(0,2)$ for all $j$, only arises when $t$ is even and $\ell=t/2$ since $\sum_{j}n_{j}=t$. In this case, for all $j$, we specialize (2.10) to determine that $c_{n_{j},m_{j}}=1$; and Lemma 5.2 implies
$\sum_{q\nmid N}T(q,n_{j},m_{j},\chi)\ =\ \frac{\sigma_{\phi}^{2}}{4}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right).$ (5.19)
Overall, our analysis supports that the main term of (5.2) is
$\frac{t!}{\left(t/2\right)!\ 2^{\ell/2}}\cdot\left(\frac{\sigma_{\phi}^{2}}{4}\right)^{t/2}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)\ =\ (t-1)!!\ \left(\frac{\sigma_{\phi}^{2}}{4}\right)^{t/2}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right).$ (5.20)
Plugging (5.20) into (3.4), we conclude that the main term of the weighted $n$th centered moment of the average one-level density in this case is
$(-2)^{n}(n-1)!!\left(\frac{\sigma_{\phi}^{2}}{4}\right)^{n/2}\ =\ (n-1)!!\ \sigma_{\phi}^{n},$ (5.21)
proving Theorem 1.1. $\Box$
@@ 5.2. Trivial Twisting Character
Although the analysis in this case is simpler than the previous case, the combinatorial arguments made to synthesize the various main term contributions in this case are more unassuming and interesting.
@@
Again, in view of Lemma 5.2, we split our consideration across two cases: $m_{j}+n_{j}\geq 3$ for some $j$, and $m_{j}+n_{j}\leq 2$ for all $j$.
@@
In the former case, consider the $j$ for which $n_{j}+m_{j}\geq 3$. For this $j$, Lemma 5.3 yields
$\sum_{q\nmid N}\widehat{\phi}\left(\frac{\log q}{\log R}\right)^{n_{j}}\frac{\log^{n_{j}}q}{q^{(n_{j}+m_{j})/2}\log^{n_{j}}R}=O\left(\frac{1}{\log^{2}R}\right)$ (5.22)
For all the other $j$, Lemma 5.2 gives
$\sum_{q\nmid N}T(q,n_{j},m_{j},\chi)\ =\ O(1).$ (5.23)
Multiplying the order of magnitude over all $j$, we find that the main term does not contribute in this case, as in (5.18).
@@
Next, we consider the case $m_{j}+n_{j}\leq 2$ for all $j$. Since $m_{j}$ and $n_{j}$ have the same parity, the set of possibilities for $(m_{j},n_{j})$ is $\{(0,2),(1,1)\}$; specializing (2.10) for these possibilities gives $c_{n_{j},m_{j}}=1$. For each $j$, based on Lemma 5.3, if $n_{j}=1$ it contributes $\phi(0)/2$ to the main term; if $n_{j}=2$ it contributes $\sigma_{\phi}^{2}/4$ to the main term. If the number of $j$’s with $n_{j}=2$ is $0\leq s\leq\lfloor t/2\rfloor$, then the number of $j$’s with $n_{j}=1$ is t2st-2s because $\sum_{j}n_{j}=t$; this case occurs $\binom{t-s}{s}$ times. Tallying the main term contribution in all these cases, we get
$\sum_{s=0}^{\lfloor t/2\rfloor}\frac{t!}{2^{s}(t-s)!}\binom{t-s}{s}\left(\frac{\phi(0)}{2}\right)^{t-2s}\left(\frac{\sigma_{\phi}^{2}}{4}\right)^{s}.$ (5.24)
Recalling (3.4), we conclude the main term of the weighted $n$th centered moment of the average one-level density is given by the following combinatorial sum:
@@
$\mathfrak{C}\ :=\ \sum_{t=0}^{n}\binom{n}{t}(-2)^{t}\sum_{s=0}^{\lfloor t/2\rfloor}\frac{t!}{2^{s}(t-s)!}\binom{t-s}{s}\left(\frac{\phi(0)}{2}\right)^{t-2s}\left(\frac{\sigma_{\phi}^{2}}{4}\right)^{s}.$ (5.25)
@@
We unravel $\mathfrak{C}$ not with the generating series trick of Soshnikov [Sos00, Lemma 2], which is characteristic in $n$-level calculations, but with an elementary, unrelated combinatorial trick of our own.
@@
$\mathfrak{C}\ =\ \phi(0)^{n}\sum_{t=0}^{n}\binom{n}{t}(-1)^{t}\sum_{s=0}^{\lfloor t/2\rfloor}\frac{t!}{(t-s)!}\binom{t-s}{s}\left(\frac{\sigma_{\phi}^{2}}{2\phi(0)^{2}}\right)^{s}.$ (5.26)
@@
$\frac{t!}{(t-s)!}\binom{t-s}{s}\left(\frac{\sigma_{\phi}^{2}}{2\phi(0)^{2}}\right)^{s}\$$=\ \frac{t!}{(t-s)!}\frac{(t-s)!}{s!(t-2s)!}\left(\frac{\sigma_{\phi}^{2}}{2\phi(0)^{2}}\right)^{s}$
$=\ \frac{t!}{s!(t-2s)!}\left(\frac{\sigma_{\phi}^{2}}{2\phi(0)^{2}}\right)^{s}$
$=\ \binom{t}{2s}\frac{(2s)!}{s!}\left(\frac{\sigma_{\phi}^{2}}{2\phi(0)^{2}}\right)^{s}.$ (5.27)
@@
$\frac{t!}{(t-s)!}\binom{t-s}{s}\left(\frac{\sigma_{\phi}^{2}}{2\phi(0)^{2}}\right)^{s}\$$=\ \binom{t}{2s}\frac{(2s)!}{s!}\left(\frac{\sigma_{\phi}^{2}}{2\phi(0)^{2}}\right)^{s}$
$=\ \binom{t}{2s}\frac{(2s)!}{2^{s}s!}\left(\frac{\sigma_{\phi}^{2}}{\phi(0)^{2}}\right)^{s}$
$=\ \binom{t}{2s}(2s-1)!!\left(\frac{\sigma_{\phi}^{2}}{\phi(0)^{2}}\right)^{s}.$ (5.28)
@@
$\frac{t!}{(t-s)!}\binom{t-s}{s}\left(\frac{\sigma_{\phi}^{2}}{2\phi(0)^{2}}\right)^{s}$ $=\ \binom{t}{2s}(2s-1)!!\left(\frac{\sigma_{\phi}^{2}}{\phi(0)^{2}}\right)^{s}$
$=\ \binom{t}{2s}\mathbb{E}\left[X^{2s}\right].$ (5.29)

### v4

@@ 5. Combinatorial Analysis
Let $\chi$ be a non-trivial character of modulus $D\geq 1$, and $\phi$ be a test function with compactly supported Fourier transform $\widehat{\phi}$. Let $n_{j}$ be a positive integer and $m_{j}$ be a non-negative integer such that $m_{j}\leq n$ and $m_{j}\equiv n_{j}\pmod{2}$. Then
$\mathcal{S}(n_{j},m_{j},\phi,\chi)\$$=\ \sum_{q\nmid N}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n_{j}}\frac{\chi(q)^{m_{j}}\log^{n_{j}}q}{q^{(n_{j}+m_{j})/2}\log^{n_{j}}Q_{k,N}}\ S_{r}\left(q^{m_{j}}\right)$
(5.9) $=\ \begin{cases}\frac{\sigma_{\phi}^{2}}{4}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)&(m_{j},n_{j})=(0,2),\\ O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)&(m_{j},n_{j})=(1,1),\\ O\left(\frac{1}{\log^{2}Q_{k,N}}\right)&\text{otherwise}.\end{cases}$
@@
First, we consider $(m_{j},n_{j})=(0,2)$. In this case, we apply Abel’s summation formula in tandem with the prime number theorem, recalling that $\phi$ is even and $\widehat{\phi}$ has compact support. Therefore, the sum under consideration equals
(5.10) $\sum_{\begin{subarray}{c}q\text{ prime};\\ q\nmid N\end{subarray}}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{2}\frac{\log^{2}q}{q\log^{2}Q_{k,N}}\ =\ \frac{\sigma_{\phi}^{2}}{4}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right).$
@@
Second, we consider $(m_{j},n_{j})=(1,1)$. We follow the same method from the previous case and appeal to Dirichlet’s theorem on primes in arithmetic progressions (the value of $\chi$ is $\pm 1$ on exactly half the primes, up to lower order terms). This gives
$\sum_{\begin{subarray}{c}q\text{ prime};\\ q\nmid N\end{subarray}}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)\frac{\chi(q)\log q}{q\log Q_{k,N}}\ S_{r}(q)$
$=$ $\ \sum_{q\text{ prime}}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)\frac{\chi(q)\log q}{q\log Q_{k,N}}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)$
$=$ $\ \sum_{q:\chi(q)=1}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)\frac{\log q}{q\log Q_{k,N}}-\sum_{q:\chi(q)=-1}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)\frac{\log q}{q\log Q_{k,N}}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)$
(5.11) $=$ $\ O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right);$
this equation is also given in [KR19, 165].
Finally, we consider $(m_{j},n_{j})\neq(0,2),(1,1)$. Since $m_{j}$ and $n_{j}$ are integers with the same parity, we know that $m_{j}+n_{j}\geq 4$ in this case, and the sum on the right-hand side of (5) converges:
$\sum_{\begin{subarray}{c}q\text{ prime};\\ q\nmid N\end{subarray}}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n_{j}}\frac{\chi(q)^{m_{j}}\log^{n_{j}}q}{q^{(m_{j}+n_{j})/2}\log^{n_{j}}Q_{k,N}}\ S_{r}\left(q^{m_{j}}\right)\ \ll_{r}$ $\ \frac{1}{\log^{n_{j}}Q_{k,N}}\sum_{q\text{ prime}}\frac{\log^{n_{j}}q}{q^{(m_{j}+n_{j})/2}}$
(5.12) $=$ $\ O\left(\frac{1}{\log^{2}Q_{k,N}}\right).$
This completes the analysis for $\chi$ non-trivial. ∎
@@
Let $\chi_{0}$ be the trivial character of modulus $D\geq 1$, and $\phi$ be a test function with compactly supported Fourier transform $\widehat{\phi}$. Let $n_{j}$ be a positive integer and $m_{j}$ be a nonnegative integer such that $m_{j}\leq n_{j}$ and $m_{j}\equiv n_{j}\pmod{2}$. Then
$\mathcal{S}(n_{j},m_{j},\phi,\chi_{0})\$$=\ \sum_{q\nmid N}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)^{n_{j}}\frac{\chi_{0}(q)^{m_{j}}\log^{n_{j}}q}{q^{(n_{j}+m_{j})/2}\log^{n_{j}}Q_{k,N}}\ S_{r}\left(q^{m_{j}}\right)$
(5.13) $=\ \begin{cases}\frac{\sigma_{\phi}^{2}}{4}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)&(m_{j},n_{j})=(0,2),\\ \frac{\phi(0)}{2}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)&(m_{j},n_{j})=(1,1),\\ O\left(\frac{1}{\log^{2}Q_{k,N}}\right)&\text{otherwise}.\end{cases}$
@@
The analysis for $(m_{j},n_{j})\neq(1,1)$ is identical to Lemma 5.2, so we need only tend to $(m_{j},n_{j})=(1,1)$. We compute the corresponding sum using the methods from the previous lemma:
$\sum_{\begin{subarray}{c}q\text{ prime};\\ q\nmid N\end{subarray}}\widehat{\phi}\left(\frac{\log q}{\log Q_{k,N}}\right)\frac{\chi_{0}(q)\log q}{q\log Q_{k,N}}\ S_{r}(q)$
[…]
This equation is also given in [KR19, 164]. ∎
With these asymptotic formulas for $\mathcal{S}(n_{j},m_{j},\phi,\chi)$, we analyze the average of interest (5.1), which we showed to be equal to (5) up to $O(\frac{\log\log(3N)}{\log Q_{k,N}})$, i.e.,
$\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(\sum\nolimits_{p\nmid N}\lambda_{\cdot}(p)\ \widehat{\phi}\left(\frac{\log p}{\log Q_{k,N}}\right)\frac{\log p}{\sqrt{p}\log Q_{k,N}}\right)^{t}\right]$
(5.15) $=\ \sum_{\ell=1}^{t}\sum_{n_{1}+\cdots+n_{\ell}=t}\sum_{\begin{subarray}{c}0\leq m_{j}\leq n_{j}\\ m_{j}\equiv n_{j}(2)\\ \forall 1\leq j\leq\ell\end{subarray}}\frac{t!}{\ell!\ n_{1}!\cdots n_{\ell}!}\left(\prod_{j=1}^{\ell}c_{n_{j},m_{j}}\cdot\mathcal{S}(n_{j},m_{j},\phi,\chi)\right)$
(5.16) $+\ O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right).$
Using our analysis of (5.15), we determine the $n$th centered moment of the one-level density. We do these separately for non-trivial and trivial $\chi$ in Sections 5.1 and 5.2, respectively.
@@ 5.1. Non-Trivial Twisting Character
Every term of (5.15) corresponds to a choice of $1\leq\ell\leq t$; a choice of $(n_{1},\dots,n_{\ell})$ such that $\sum_{j=1}^{\ell}n_{j}=t$; and a choice of $(m_{1},\dots,m_{\ell})$ such that $m_{j}\equiv n_{j}\pmod{2}$ and $0\leq m_{j}\leq n_{j}$ for all $j$. In view of Lemma 5.2, we split our consideration of these terms into two cases: terms for which $(m_{j},n_{j})\neq(0,2)$ for some $j$, and the term for which $(m_{j},n_{j})=(0,2)$ for all $j$.
For terms of the former type, consider the $j$ for which $(m_{j},n_{j})\neq(0,2)$. For this $j$, Lemma 5.2 yields
(5.17) $\mathcal{S}(n_{j},m_{j},\phi,\chi)\ =\ O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right);$
and for all the other $j$,
(5.18) $\mathcal{S}(n_{j},m_{j},\phi,\chi)\ =\ O(1).$
Multiplying the order of magnitude over all $j$, each of these terms contributes
(5.19) $\frac{t!}{\ell!\ n_{1}!\cdots n_{\ell}!}\left(\prod_{j=1}^{\ell}c_{n_{j},m_{j}}\cdot\mathcal{S}(n_{j},m_{j},\phi,\chi)\right)\ =\ O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right).$
Since the number of terms of this type depends only on $t$, the total contribution is also $O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)$, the implied constant depending on $t$, $r$, and $D$.
As for the latter case, we consider the term for which $(m_{j},n_{j})=(0,2)$ for all $j$. This arises only when $t$ is even and $\ell=t/2$ because $\sum_{j=1}^{\ell}n_{j}=t$. It follows from (2) that $c_{n_{j},m_{j}}=1$ for all $j$; and Lemma 5.2 gives
(5.20) $\mathcal{S}(n_{j},m_{j},\phi,\chi)\ =\ \frac{\sigma_{\phi}^{2}}{4}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)$
for all $j$. Therefore, the contribution of this term to (5.15) is
(5.21) $\frac{t!}{\left(t/2\right)!\ 2^{\ell/2}}\cdot\left(\frac{\sigma_{\phi}^{2}}{4}\right)^{t/2}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)\ =\ (t-1)!!\ \left(\frac{\sigma_{\phi}^{2}}{4}\right)^{t/2}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right).$
Adding the contributions of all these terms (5.19) and (5.21), we have that
$\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(\sum\nolimits_{p\nmid N}\lambda_{\cdot}(p)\ \widehat{\phi}\left(\frac{\log p}{\log Q_{k,N}}\right)\frac{\log p}{\sqrt{p}\log Q_{k,N}}\right)^{t}\right]$
(5.22) $=\ (t-1)!!\ \left(\frac{\sigma_{\phi}^{2}}{4}\right)^{t/2}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)$
Finally, taking $t=n$ in (5.22), we deduce that the $n$th centered moment of the $w_{\chi_{0},r}$-weighted one-level density of $\mathcal{F}_{k}(N)^{\prime}$, originally formulated in (3.5), is
(5.23) $\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(-2\sum_{p\nmid N}\lambda_{\cdot}(p)\widehat{\phi}\left(\frac{\log p}{\log Q_{k,N}}\right)\frac{\log p}{\sqrt{p}\log Q_{k,N}}+O\left(\frac{\log\log(kN)}{\log Q_{k,N}}\right)\right)^{n}\right]$
(5.24) $=\ (-2)^{n}\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(\sum_{p\nmid N}\lambda_{\cdot}(p)\widehat{\phi}\left(\frac{\log p}{\log Q_{k,N}}\right)\frac{\log p}{\sqrt{p}\log Q_{k,N}}\right)^{n}\right]+O\left(\frac{\log\log(kN)}{\log Q_{k,N}}\right)$
(5.25) $=\ (-2)^{n}\ (n-1)!!\ \left(\frac{\sigma_{\phi}^{2}}{4}\right)^{n/2}+O\left(\frac{\log\log(kN)}{\log Q_{k,N}}\right)$
(5.26) $=\ (n-1)!!\ \sigma_{\phi}^{n}+O\left(\frac{\log\log(kN)}{\log Q_{k,N}}\right),$
proving Theorem 1.1 in the case that $\chi$ is non-trivial.
@@ 5.2. Trivial Twisting Character
Although the analysis for trivial $\chi=\chi_{0}$, the combinatorial arguments made to synthesize the various term-wise contributions in this case are more unassuming.
@@
In view of Lemma 5.3, we split our consideration of the terms of (5.15) into two cases: terms for which $m_{j}+n_{j}\geq 3$ for some $j$, and terms for which $m_{j}+n_{j}\leq 2$ for all $j$.
@@
For terms of the the former type, consider the $j$ for which $n_{j}+m_{j}\geq 3$. For this $j$, Lemma 5.3 yields
(5.27) $\mathcal{S}(n_{j},m_{j},\phi,\chi_{0})\ =\ O\left(\frac{1}{\log^{2}Q_{k,N}}\right);$
and for all the other $j$,
(5.28) $\mathcal{S}(n_{j},m_{j},\phi,\chi_{0})\ =\ O(1).$
Multiplying the order of magnitude over all $j$, we find that each of these terms contributes
(5.29) $\frac{t!}{\ell!\ n_{1}!\cdots n_{\ell}!}\left(\prod_{j=1}^{\ell}c_{n_{j},m_{j}}\cdot\mathcal{S}(n_{j},m_{j},\phi,\chi_{0})\right)\ =\ O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right).$
@@
Next, we consider the terms for which $m_{j}+n_{j}\leq 2$ for all $j$. Since $m_{j}$ and $n_{j}$ have the same parity, the set of possibilities for $(m_{j},n_{j})$ is $\{(0,2),(1,1)\}$; specializing (2) for these possibilities gives $c_{n_{j},m_{j}}=1$ for all $j$. Based on Lemma 5.3, if $n_{j}=1$ it contributes $\phi(0)/2$ to the main term; if $n_{j}=2$ it contributes $\sigma_{\phi}^{2}/4$ to the main term. If the number of $j$’s with $n_{j}=2$ is $0\leq s\leq\lfloor t/2\rfloor$, then the number of $j$’s with $n_{j}=1$ is t2st-2s because $\sum_{j}n_{j}=t$; this case occurs $\binom{t-s}{s}$ times. Tallying the contribution in all these cases, we get
$\frac{t!}{\ell!\ n_{1}!\cdots n_{\ell}!}\left(\prod_{j=1}^{\ell}c_{n_{j},m_{j}}\cdot\mathcal{S}(n_{j},m_{j},\phi,\chi_{0})\right)$
(5.30) $=\ \sum_{s=0}^{\lfloor t/2\rfloor}\frac{t!}{2^{s}(t-s)!}\binom{t-s}{s}\left(\frac{\phi(0)}{2}\right)^{t-2s}\left(\frac{\sigma_{\phi}^{2}}{4}\right)^{s}.$
Overall, adding (5.29) and (5.2), we have that
$\mathcal{A}^{w_{\chi_{0},r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(\sum\nolimits_{p\nmid N}\lambda_{\cdot}(p)\ \widehat{\phi}\left(\frac{\log p}{\log Q_{k,N}}\right)\frac{\log p}{\sqrt{p}\log Q_{k,N}}\right)^{t}\right]$
(5.31) $=\ \sum_{s=0}^{\lfloor t/2\rfloor}\frac{t!}{2^{s}(t-s)!}\binom{t-s}{s}\left(\frac{\phi(0)}{2}\right)^{t-2s}\left(\frac{\sigma_{\phi}^{2}}{4}\right)^{s}+O\left(\frac{\log\log(3N)}{\log Q_{k,N}}\right)$
Using (5.31), we now calculate the $n$th centered moment of the $w_{\chi_{0},r}$-weighted one-level density of $\mathcal{F}_{k}(N)^{\prime}$, originally formulated in (3.5):
$\mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(\phi(0)-2\sum_{p\nmid N}\lambda_{\cdot}(p)\widehat{\phi}\left(\frac{\log p}{\log Q_{k,N}}\right)\frac{\log p}{\sqrt{p}\log Q_{k,N}}+O\left(\frac{\log\log(kN)}{\log Q_{k,N}}\right)\right)^{n}\right]$
$=\ \sum_{t=0}^{n}\binom{n}{t}\phi(0)^{n-t}(-2)^{t}\ \mathcal{A}^{w_{\chi,r}}_{\mathcal{F}_{k}(N)^{\prime}}\left[\left(\sum_{p\nmid N}\lambda_{\cdot}(p)\widehat{\phi}\left(\frac{\log p}{\log(k^{2}N)}\right)\frac{\log p}{\sqrt{p}\log Q_{k,N}}\right)^{t}\right]$
$+\ O\left(\frac{\log\log(kN)}{\log Q_{k,N}}\right)$
$=\ \sum_{t=0}^{n}\binom{n}{t}\phi(0)^{n-t}(-2)^{t}\sum_{s=0}^{\lfloor t/2\rfloor}\frac{t!}{2^{s}(t-s)!}\binom{t-s}{s}\left(\frac{\phi(0)}{2}\right)^{t-2s}\left(\frac{\sigma_{\phi}^{2}}{4}\right)^{s}$
(5.32) $+\ O\left(\frac{\log\log(kN)}{\log Q_{k,N}}\right).$
We denote the main term above by
@@
(5.33) $\mathfrak{C}\ :=\ \sum_{t=0}^{n}\binom{n}{t}\phi(0)^{n-t}(-2)^{t}\sum_{s=0}^{\lfloor t/2\rfloor}\frac{t!}{2^{s}(t-s)!}\binom{t-s}{s}\left(\frac{\phi(0)}{2}\right)^{t-2s}\left(\frac{\sigma_{\phi}^{2}}{4}\right)^{s}.$
@@
We unravel $\mathfrak{C}$ not with the generating series trick of Soshnikov [Sos00, Lemma 2], which is characteristic in $n$-level calculations, but with an elementary combinatorial trick of our own.
@@
(5.34) $\mathfrak{C}\ =\ \phi(0)^{n}\sum_{t=0}^{n}\binom{n}{t}(-1)^{t}\sum_{s=0}^{\lfloor t/2\rfloor}\frac{t!}{(t-s)!}\binom{t-s}{s}\left(\frac{\sigma_{\phi}^{2}}{2\phi(0)^{2}}\right)^{s}.$
@@
(5.35) $\frac{t!}{(t-s)!}\binom{t-s}{s}\left(\frac{\sigma_{\phi}^{2}}{2\phi(0)^{2}}\right)^{s}\$$=\ \binom{t}{2s}\frac{(2s)!}{s!}\left(\frac{\sigma_{\phi}^{2}}{2\phi(0)^{2}}\right)^{s}.$
@@
(5.36) $\frac{t!}{(t-s)!}\binom{t-s}{s}\left(\frac{\sigma_{\phi}^{2}}{2\phi(0)^{2}}\right)^{s}\$$=\ \binom{t}{2s}(2s-1)!!\left(\frac{\sigma_{\phi}^{2}}{\phi(0)^{2}}\right)^{s}.$
@@
(5.37) $\frac{t!}{(t-s)!}\binom{t-s}{s}\left(\frac{\sigma_{\phi}^{2}}{2\phi(0)^{2}}\right)^{s}$ $=\ \binom{t}{2s}\mathbb{E}\left[X^{2s}\right].$